Lynn can mow the lawn in 3 hours using a riding lawn mower. Her brother can mow the same lawn in 7 hours using a push mower. How long would it take them to mow the lawn working together?
Answer by venugopalramana(3286)

SEE THE FOLLOWING EXAMPLES AND TRY.
------------------------------------------------------
Rate-of-work-word-problems/30847: Please help me with this problem. I think it needs to be set up as a linear equation. Also, it would help if you could explain it step by step for me. Thanks so much! ** An empty swimming pool can be filled in 10 hours. When full, the pool can be drained in 16 hours. How long will it take to fill the pool if the drain is left open?
1 solutions
Answer 17587 by venugopalramana(1167)on 2006-03-20 21:52:19 
Please help me with this one. ** An empty swimming pool can be filled in 10 hours.
SO IN 1 HOUR WE CAN FILL 1/10 POOL
When full, the pool can be drained in 16 hours.
SO IN 1 HOUR 1/16 POOL WILL BE DRAINED
How long will it take to fill the pool if the drain is left open?
SO IN 1 HOUR 1/10 - 1/16 = (8-5)/80 = 3/80 POOL WILL BE FILLED
SO IT TAKES 80/3 = 26.67 HRS TO FILL THE POOL

Miscellaneous_Word_Problems/30717: Please help me with this one. ** An empty swimming pool can be filled in 10 hours. When full, the pool can be drained in 16 hours. How long will it take to fill the pool if the drain is left open?
1 solutions
Answer 17442 by venugopalramana(1167)on 2006-03-19 21:28:52 
Please help me with this one. ** An empty swimming pool can be filled in 10 hours.
SO IN 1 HOUR WE CAN FILL 1/10 POOL
When full, the pool can be drained in 16 hours.
SO IN 1 HOUR 1/16 POOL WILL BE DRAINED
How long will it take to fill the pool if the drain is left open?
SO IN 1 HOUR 1/10 - 1/16 = (8-5)/80 = 3/80 POOL WILL BE FILLED
SO IT TAKES 80/3 = 26.67 HRS TO FILL THE POOL

Mixture_Word_Problems/30424: Dennis can eat a pie in 6 min. David can eat a pie in 12 min. Robert can eat a pie in 14.639145073927682623334797247841 min. If they share the same pie, how lomg will it take the three guys to scarf down the whole thing. (Hint: First figure out how many pies each guy can eat in an hour)
1 solutions
Answer 17065 by venugopalramana(1167)on 2006-03-15 22:41:46 
Dennis can eat a pie in 6 min. David can eat a pie in 12 min. Robert can eat a pie in 14.639145073927682623334797247841 min. If they share the same pie, how lomg will it take the three guys to scarf down the whole thing. (Hint: First figure out how many pies each guy can eat in an hour)
DENIS CAN EAT IN 1 MINUTE....1/6 =0.166666667 PIE
DAVID CAN EAT IN 1 MINUTE....1/12 =0.083333333 PIE
ROBERT CAN EAT IN 1 MINUTE...1/14.63914507=0.06831 PIE
HENCE THE 3 TOGETHER CAN EAT IB 1 MINUTE=0.31831 PIE
HENCE THEY TAKE ...1/0.31831 = 3.14159153 MINUTES TO EAT THE PIE.

Rate-of-work-word-problems/29587: need help plz...heres the question "There are three friends named Allan, Bobby and Charlie. The three friends want to know their individual rate in finishing a job. Allan and Bobby can finish the job in 42 days, Bobby and Charlie can finish the job in 31 days, and Allan and Charlie can finish the job in 20 days. Solve the rate of each individual."
1 solutions
Answer 16715 by venugopalramana(1167)on 2006-03-12 11:04:12 
need help plz...heres the question "There are three friends named Allan, Bobby and Charlie. The three friends want to know their individual rate in finishing a job. Allan and Bobby can finish the job in 42 days, Bobby and Charlie can finish the job in 31 days, and Allan and Charlie can finish the job in 20 days. Solve the rate of each individual."
LET ALLAN TAKE A DAYS TO DO THE JOB ALONE
HENCE ALLAN ALONE CAN DO IN 1 DAY 1/A JOB
LET BOBBY TAKE B DAYS TO DO THE JOB ALONE
HENCE BOBBY ALONE CAN DO IN 1 DAY 1/B JOB
LET CHARLIE TAKE C DAYS TO DO THE JOB ALONE
HENCE CHARLIE ALONE CAN DO IN 1 DAY 1/C JOB
FROM ABOVE WE GET ...
ALLAN AND BOBBY CAN DO IN 1 DAY ..1/A +1/B =(A+B)/AB JOB..
SO DAYS THEY TAKE TO COMPLETE THE JOB =AB/(A+B)=42..OR..1/A+1/B=1/42..........I BOBBY AND CHARLIE CAN DO IN 1 DAY ..1/B +1/C =(B+C)/BC JOB..
SO DAYS THEY TAKE TO COMPLETE THE JOB =BC/(B+C)=31..OR..1/B+1/C=1/31.........II
ALLAN AND CHARLIE CAN DO IN 1 DAY ..1/A +1/C =(A+C)/AC JOB..
SO DAYS THEY TAKE TO COMPLETE THE JOB =AC/(A+C)=20..OR..1/A+1/C=1/20........III
EQNI+EQNII+EQNIII GIVES...
2{(1/A)+(1/B)+(1/C)}=(1/42)+(1/31)+(1/20)
(1/A)+(1/B)+(1/C)=(1/2)*{(1/42)+(1/31)+(1/20)}.........IV
EQN.IV-EQN.I GIVES
1/C= (1/2)*{(1/42)+(1/31)+(1/20)}-1/42....OR.....C=34.2 DAYS
EQN.IV-EQN.II GIVES.......
1/A=(1/2)*{(1/42)+(1/31)+(1/20)}- 1/31...OR......A=48.1 DAYS
EQN.IV-EQN.III GIVES....
1/B=(1/2)*{(1/42)+(1/31)+(1/20)}- 1/20....OR.....B=329.6 DAYS